# Spontaneous emission and lasing in photonic time crystals

### v1

@@ Abstract
To quantitatively capture the evolving lineshape within the momentum gap, we analyze the behavior of the key parameters in Eq.(3): the real and imaginary parts of the $I_{\alpha}$, which control the weights of the symmetric and anti-symmetric components, and the mode linewidth $\Gamma_{\alpha}$. Both quantities are extracted by fitting the measured spectra to the two-mode expression. Figure3 e shows that $|\Im[I_{\alpha}]|$ decreases with increasing modulation frequency, as the mode approaches the center of the gap, consistent with the transition from a skewed to a symmetric spectral profile. In contrast, $\Re[I_{\alpha}]$ remains nearly constant throughout. Among the two gap modes, the one with the larger linewidth, $\Gamma_{+}$, is strongly damped and buried beneath the noise floor, making reliable extraction of its parameters difficult. We therefore focus on the less-damped mode, whose linewidth is $\Gamma_{-}$, as its spectral contribution remains clearly visible across the modulation range. As the modulation frequency increases, the gap mode moves closer to the center of the momentum gap, where the imaginary part of its eigenfrequency—and thus the linewidth—tends to decrease (Fig. 3f). This trend is consistent with the characteristic band structure of a PTC, where the gap modes exhibit minimal damping near the gap center and stronger attenuation toward the edges.
Once the modulation strength exceeds the point where parametric gain overcomes the intrinsic and radiative losses of the finite-sized PTC, it enters an oscillating regime (Fig. 4): a narrow-band oscillation appears (so called PTC laser) and eventually saturates because of the varactor nonlinearity in each unit cell. In contrast to conventional lasers that rely on stimulated emission, the PTC laser achieves light amplification through parametric gain induced by periodic modulation, without requiring a population-inverted gain medium. The term “laser” is used here in a generalized sense, referring to the emergence of narrow-band, oscillating radiation above a well-defined threshold. Here, the threshold power can be understood as the point at which the mode inside the momentum gap begins to grow in time, i.e., when the imaginary part of quasi-eigenfrequency crosses zero, $\Gamma_{-}=0$. More precisely, the threshold is reached at the critical modulation depth $\delta_{c}=4\omega_{k}^{-1}\sqrt{\Delta^{2}+\gamma^{2}}$, where $\Delta\equiv\Omega/2-\omega_{k}$ is the detuning from exact parametric resonance and $\gamma$ is the intrinsic damping rate of the resonator Bloch mode (see Supplementary Material B for details). In this strongly driven regime, the linear Floquet model breaks down, so the earlier LDOS analysis no longer applies. In this regime, the system behaves as a one-dimensional array of nonlinear LC resonators with capacitances modulated at frequency $\Omega$. Such systems are known to host an activated discrete-time-crystalline phase, characterized by synchronized sub-harmonic oscillations that spontaneously break the discrete time-translation symmetry of the modulation [36]. Time-crystalline phases have attracted growing interest across both classical and quantum platforms [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
We have shown that a PTC, implemented here with an array of varactor-loaded microwave resonators, can tailor the electromagnetic vacuum as effectively as spatial structuring. By measuring the radiated power spectral density seeded by noise in the unit cell circuits, we directly mapped the spectrally resolved LDOS and observed it reshape in real time as we tuned the drive frequency and depth. The measured lineshape, its skewness, and the steep growth of the peak all track the non-Hermitian Floquet theory, including the Petermann enhancement that accompanies the EPs at a momentum gap edge. These experiments therefore place the LDOS, arguably the most fundamental quantity in light–matter interaction, squarely under temporal control. The technique paves the way to dynamic Purcell engineering, tests of dynamical Casimir analogues, and on-demand switching between emission, absorption and even spontaneous excitation regimes in quantum emitters coupled to PTCs.
@@
Supplemental Material: Spontaneous emission and lasing in photonic time crystals
@@
A Circuit model derivation of microwave time crystal Hamiltonian
We model the microwave time crystal as a one-dimensional array of LC resonators side-coupled to a bus waveguide (Fig. 1a of the main text), as in our previous work[51]. The Lagrangian is
@@ A Circuit model derivation of microwave time crystal Hamiltonian
Here, $Q_{n}^{X}$ is the charge on element $X$ in unit cell $n$, and $\dot{Q}_{n}^{X}$ is the corresponding current. The lumped parameters $L_{0}$, $L_{s}$ and $C_{0}$ specify, respectively, the series inductance, shunt inductance, and ground capacitance that comprise the discretized waveguide. Each resonator contains an inductance $L_{r}$ and a time-dependent capacitance $C_{r}(t)$. The term $M$ denotes the mutual inductance between neighboring resonators, with only nearest-neighbor coupling assumed. The parameter $C_{g}$ represents the coupling capacitance between the resonator and the waveguide. Moving to momentum space, we label the Bloch modes by an integer $m=0,1,\dots,N-1$ and define the discrete wavenumbers $k_{m}\equiv 2\pi m/N$. The discrete Fourier transform is then
@@
where the coupling strength between the waveguide and the resonators is given by $g_{k}=\sqrt{C_{0}C_{c}\Omega_{k}\omega_{k}}/(2C_{g})$
B Floquet Hamiltonian of microwave time crystals
@@
The dominant intrinsic loss arises from Ohmic dissipation in the resonators. We model it by inserting a series resistor $R$ into each resonator (Fig. 1a of the main text) and define the Rayleigh dissipation function in terms of the resonator currents:
@@
$=\sum_{k}\mathbf{x}_{k}^{\dagger}\hat{F}_{k}\mathbf{x}_{k},$
@@
$(\omega_{\alpha}+q\Omega)\psi_{k,\alpha}^{q}=\sum_{r}\mathbf{H}_{k}^{q-r}\psi_{k,\alpha}^{r},$ (S19)
@@
and the corresponding Floquet eigenvector is $\mathbf{F}_{k,\alpha}=[\cdots;\psi_{k,\alpha}^{-1};\psi_{k,\alpha}^{0};\psi_{k,\alpha}^{1};\cdots]^{T}$. The off-diagonal blocks $\mathbf{H}_{k}^{m}\;(m\neq 0)$ mediate the coupling between the original band, spanned by the eigenmodes of $\mathbf{H}_{k}^{0}$, and its frequency-shifted replicas, spanned by the eigenmodes of $\mathbf{H}_{k}^{0}+m\Omega I_{4}$. The eigenvalues $\omega_{\alpha}$ obtained from $\mathbf{H}_{k}^{F}$ therefore constitute the Floquet band structure of the microwave time crystal.
When the driving frequency is close to twice the resonator Bloch mode frequency, $\Omega\simeq 2\omega_{k}$, only the positive‑frequency at $\omega_{k}$ band and the Floquet replica of its negative-frequency counterpart at $\Omega-\omega_{k}$ remain near resonance; all other sidebands are far detuned. In this scenario, the coupling between these two bands which gives rise to the formation of the momentum gap can be described by the following reduced $\times 2$ 2 non-Hermitian Floquet Hamiltonian:
@@
where $\Delta\equiv\frac{\Omega}{2}-\omega_{k}$ is the detuning from exact resonance. For $\delta<\delta_{c}$ both modes remain lossy. When $\delta>\delta_{c}$ one mode acquires net positive gain, signaling the onset of a parametric instability, or PTC-lasing transition.
C Power dissipation and local density of states in microwave time crystals
@@
All other terms in the Lagrangian remain as in Eq. S1. Here, the applied voltage $V_{d}(t)$ drives a localized, oscillating dipole, i.e., the resonator charge $Q_{n_{0}}^{L_{r}}$ at site $n_{0}$. Within linear response, the time-averaged power delivered by thisdipole is proportional to the local density of states [34]. Applying the inverse discrete Fourier transform to $Q_{n_{0}}^{L_{r}}$, the drive term becomes
@@
[…]
@@
such that the total LDOS satisfies $\rho(\omega;n_{0})=N^{-1}\sum_{k}\rho_{k}(\omega)$. Combining Eqs.Eqs. S35 and S36 , we obtain a compact expression that relates the dissipated power to the LDOS:

### v2

@@ Abstract
To quantitatively capture the evolving lineshape within the momentum gap, we analyze the behavior of the key parameters in Eq. (3): the real and imaginary parts of $I_{\alpha}$, which control the weights of the symmetric and antisymmetric components, and the mode linewidth $\Gamma_{\alpha}$. These quantities are extracted by fitting the measured spectra to the two-mode expression. Figure 3e shows that $|\Im[I_{\alpha}]|$ decreases with increasing modulation frequency, as the mode approaches the center of the gap, consistent with the transition from a skewed to a symmetric spectral profile. In contrast, $\Re[I_{\alpha}]$ remains nearly constant throughout. Among the two gap modes, the one with the larger linewidth, $\Gamma_{+}$, is strongly damped and buried beneath the noise floor, making reliable extraction of its parameters difficult. We therefore focus on the less-damped mode, whose linewidth is $\Gamma_{-}$, as its spectral contribution remains clearly visible across the modulation range. As the modulation frequency increases, the gap mode moves closer to the center of the momentum gap, where the imaginary part of its eigenfrequency—and thus the linewidth—tends to decrease (Fig. 3f). This trend is consistent with the characteristic band structure of a PTC, where the gap modes exhibit minimal damping near the gap center and stronger attenuation toward the edges.
Once the modulation-induced parametric gain overcomes the losses of the finite-sized PTC, the in-gap Floquet mode becomes unstable and the system enters a self-oscillating (parametric-oscillator) regime (Fig. 4). Above threshold, a coherent narrow-band oscillation appears; we call this regime a “PTC laser” to emphasize its laser-like coherence despite its parametric origin, and it saturates due to varactor nonlinearities. Unlike conventional population-inversion lasers, the amplification here is supplied by the external time-periodic modulation and is therefore best viewed as modulation-induced parametric gain rather than material gain. We therefore interpret the oscillation as modulation-induced parametric self-oscillation, closely analogous to a degenerate optical parametric oscillator [eckardt1991optical]. The observed threshold is consistent with the onset of a parametric instability of the in-gap mode, i.e., when the imaginary part of its quasi-eigenfrequency approaches zero ($\Gamma_{-}\to 0$) within the linear Floquet model. Within a near-resonant, two-mode Floquet approximation, the instability occurs when the modulation depth reaches the critical value
$\delta_{c}=\frac{4}{\omega_{k}}\sqrt{\Delta^{2}+\gamma^{2}},$ (6)
where $\Delta\equiv\Omega/2-\omega_{k}$ is the detuning from exact parametric resonance and $\gamma$ denotes the damping rate of the resonator Bloch mode (see Supplementary Material B for details). Beyond threshold, the oscillation amplitude grows to a point where the linear Floquet description breaks down: the nonlinear response of the varactors becomes essential, rendering the earlier LDOS analysis inapplicable. In this strongly driven regime, the PTC behaves effectively as a one-dimensional array of nonlinear, parametrically driven LC oscillators. From the viewpoint of coupled nonlinear parametric oscillators, such systems are natural candidates for activated discrete-time-crystalline behavior, exhibiting synchronized subharmonic oscillations that spontaneously break the discrete time-translation symmetry of the drive [yao2020classical]. While distinct from the photonic time crystal considered here, discrete time-crystalline phases have attracted wide interest across classical and quantum platforms [PhysRevLett.123.124301, pizzi2021aclassical, PhysRevE.100.060105, taheri2022all, giergiel2023discrete, PhysRevLett.127.140603, yi2024theory, choi2017observation, zhang2017observation, ho2017critical, rovny2018p, liu2023photonic, kyprianidis2021observation, mi2022time].
@@
We have shown that a PTC, implemented here with an array of varactor‑loaded microwave resonators, can tailor the electromagnetic vacuum in a manner analogous to spatial structuring, but using time as the control knob. By measuring the noise-seeded radiated power spectral density emitted from the unit-cell circuits, we directly mapped the spectrally resolved LDOS and observed it reshape as we varied the drive frequency and modulation depth. The measured lineshape, its skewness, and the steep growth of the peak all follow the predictions of non‑Hermitian Floquet theory, including the Petermann factor enhancement associated with exceptional points at the momentum‑gap edges. These experiments therefore place the LDOS—arguably the most fundamental quantity in light–matter interaction—squarely under active temporal control. Looking ahead, this approach opens a route to dynamic Purcell engineering, tabletop tests of dynamical Casimir analogs, and, in suitably coupled platforms, on‑demand switching between emission, absorption, and even spontaneous excitation regimes for quantum emitters interfaced with photonic time crystals.
@@
Supplemental Material: Analogs of spontaneous emission and lasing in photonic time crystals
@@
A Circuit model derivation of photonic time crystal Hamiltonian
We model the photonic time crystal as a one-dimensional array of LC resonators side-coupled to a bus waveguide (Fig. 1a of the main text), as in our previous work[park2022revealing]. The Lagrangian is
@@ A Circuit model derivation of photonic time crystal Hamiltonian
Here, $Q_{n}^{X}$ denotes the charge on element $X$ in unit cell $n$, and $\dot{Q}_{n}^{X}$ is the corresponding current. The lumped parameters $L_{0}$, $L_{s}$ and $C_{0}$ specify, respectively, the series inductance, shunt inductance, and ground capacitance that comprise the discretized waveguide. Each resonator contains an inductance $L_{r}$ and a time-dependent capacitance $C_{r}(t)$. The term $M$ denotes the mutual inductance between neighboring resonators, with only nearest-neighbor coupling assumed. The parameter $C_{g}$ represents the coupling capacitance between the resonator and the waveguide. Moving to momentum space, we label the Bloch modes by an integer $m=0,1,\dots,N-1$ and define the discrete wavenumbers $k_{m}\equiv 2\pi m/N$. The discrete Fourier transform is then
@@
where the coupling strength between the waveguide and the resonators is given by $g_{k}=\sqrt{C_{0}C_{c}\Omega_{k}\omega_{k}}/(2C_{g})$ .
B Floquet Hamiltonian of photonic time crystals
@@
The dominant intrinsic loss arises from Ohmic dissipation in the resonators. We model it by inserting a resistor $R$ into each resonator (Fig. 1a of the main text) and define the Rayleigh dissipation function in terms of the resonator currents:
@@
$=\sum_{k}\mathbf{x}_{k}^{\dagger}\hat{D}_{k}\mathbf{x}_{k},$
@@
$(\omega_{\alpha}+q\Omega)\psi_{k,\alpha}^{q}=\sum_{r}\mathbf{H}_{k}^{q-r}\psi_{k,\alpha}^{r},\$(S19)
@@
and the corresponding Floquet eigenvector is $\mathbf{F}_{k,\alpha}=[\cdots;\psi_{k,\alpha}^{-1};\psi_{k,\alpha}^{0};\psi_{k,\alpha}^{1};\cdots]^{T}$. The off-diagonal blocks $\mathbf{H}_{k}^{m}\;(m\neq 0)$ mediate the coupling between the original band, spanned by the eigenmodes of $\mathbf{H}_{k}^{0}$, and its frequency-shifted replicas, spanned by the eigenmodes of $\mathbf{H}_{k}^{0}+m\Omega I_{4}$. The eigenvalues $\omega_{\alpha}$ obtained from $\mathbf{H}_{k}^{F}$ therefore constitute the Floquet band structure of the photonic time crystal.
When the driving frequency is close to twice the resonator Bloch mode frequency, $\Omega\simeq 2\omega_{k}$, only the positive‑frequency at $\omega_{k}$ band and the Floquet replica of its negative-frequency counterpart at $\Omega-\omega_{k}$ remain near resonance; all other sidebands are far detuned. In this scenario, the coupling between these two bands which gives rise to the formation of the momentum gap can be described by the following reduced $2\times 2$ non-Hermitian Floquet Hamiltonian:
@@
where $\Delta\equiv\frac{\Omega}{2}-\omega_{k}$ is the detuning from exact resonance. For $\delta<\delta_{c}$ both modes remain lossy. When $\delta>\delta_{c}$ one mode acquires net positive gain, signaling the onset of a parametric instability, or PTC-laser (self-oscillation) transition.
C Power dissipation and local density of states in photonic time crystals
@@
All other terms in the Lagrangian remain as in Eq. S1. Here, the applied voltage $V_{d}(t)$ drives a localized, oscillating dipole, i.e., the resonator charge $Q_{n_{0}}^{L_{r}}$ at site $n_{0}$. Within linear response, the time-averaged power delivered by this dipole is proportional to the local density of states [park2025spontaneous_SI]. Applying the inverse discrete Fourier transform to $Q_{n_{0}}^{L_{r}}$, the drive term becomes
@@
$f_{k}=\sqrt{\frac{C_{c}\omega_{k}}{2N}}\,e^{ikn_{0}},\quad\mathbf{s}=[\cdots;\,\mathbf{0};\,[0,-1,0,1]^{T};\,\mathbf{0};\,\cdots],$ (S31)
@@
such that the total LDOS satisfies $\rho(\omega;n_{0})=N^{-1}\sum_{k}\rho_{k}(\omega)$. Combining Eqs. S35 and S36 , we obtain a compact expression that relates the dissipated power to the LDOS:
@@
D Experimental Implementation of the PTCs
Figure D.1: (a) Schematic of the one-dimensional array of 12 coupled LC resonators embedded in a waveguide. (b) Schematic layout and photograph of the fabricated unit cell, incorporating a varactor diode for time-periodic capacitance modulation via a DC-biased AC voltage. The unit cell is designed with the following geometrical parameters: $\mathrm{h}_{1}=40\ \mathrm{mm}$, $\mathrm{h}_{2}=1\ \mathrm{mm}$, $\mathrm{h}_{3}=3\ \mathrm{mm}$, $\mathrm{h}_{4}=10\ \mathrm{mm}$, $\mathrm{w}_{1}=105\ \mathrm{mm}$ and $\mathrm{w}_{2}=27.5\ \mathrm{mm}$. (c) Schematic of the experimental setup.
To implement the unit cell of the LC resonator, a metallic split-ring resonator (SRR) was patterned onto a printed circuit board (Fig. D.1b). Each unit cell incorporates a varactor diode (SMV1247, Skyworks), whose capacitance can be modulated by applying a DC-biased AC voltage across it. As a result, the resonant characteristics of the unit cell undergo time-periodic modulation. In the experiment, a DC bias voltage of 2.6 V was applied. To construct a one-dimensional array of coupled LC resonators, 12 unit cells were placed inside a waveguide (Fig. D.1a). The unit cells were arranged with a center-to-center spacing of 10 mm and the waveguide was terminated at both ends with 50 $\Omega$ loads to suppress reflections. The complete experimental setup is shown in Fig. D.1c.
The modulation signal (AC voltage) is generated by a signal generator (N5171B, Keysight) and amplified by a radio-frequency amplifier (RUM43020-10, RFHIC). The amplified AC voltage is then DC-biased using a bias tee and subsequently distributed through a two-way splitter, followed by a six-way splitter. Additional two-way splitters are inserted just before the resonator array, with each splitter enabling individual signal measurement from a corresponding unit cell. To prevent unwanted DC current flow, a DC block is inserted at one end of the two-way splitter. Additionally, a low-pass filter (LPF) is placed before the oscilloscope (DPO 70804, Tektronix) or spectrum analyzer (E4404B, Agilent) to suppress reflections of the modulation signal.
To examine the predicted enhancement of the LDOS near the momentum-gap frequency, we operate the structure with no coherent input; the only drive is broadband white noise arising from Johnson–Nyquist thermal fluctuations in combination with residual technical noise. These fluctuations act as spatially distributed, broadband noise sources that probe the linear response of the PTC. The measured noise-driven radiated spectrum closely follows the LDOS lineshape derived from the non-Hermitian Floquet framework, providing an experimental proxy for the frequency dependence of the spontaneous-emission rate in the weak-coupling limit.